# Tomoya Shiota,∗,†,a,b Kenji Ishihara,†,b and Wataru Mizukami∗,a,b

### v2

@@ Acknowledgements
We thank Nobuki Inoue, Tuan Minh Do, and Ryo Watanabe for fruitful discussions. This project was supported by funding from the MEXT Quantum Leap Flagship Program (MEXTQLEAP) through Grant No. JPMXS0120319794, and the JST COI-NEXT Program through Grant No. JPMJPF2014. The completion of this research was partially facilitated by the JSPS Grants-in-Aid for Scientific Research (KAKENHI), specifically Grant Nos. JP23H03819 and JP21K18933. We thank the Supercomputer Center, the Institute for So1lid State Physics, the University of Tokyo, for allowing us to use these facilities. This work was also achieved using SQUID at the Cybermedia Center, Osaka University.
@@
  • 85 https://github.com/Qulacs-Osaka/scikit-qulacs.

### v3

@@ Acknowledgements
We thank Nobuki Inoue, Tuan Minh Do, and Ryo Watanabe for fruitful discussions. This project was supported by funding from the MEXT Quantum Leap Flagship Program (MEXTQLEAP) through Grant No. JPMXS0120319794, and the JST COI-NEXT Program through Grant No. JPMJPF2014. The completion of this research was partially facilitated by the JSPS Grants-in-Aid for Scientific Research (KAKENHI), specifically Grant Nos. JP23H03819 and JP21K18933. We thank the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo, for allowing us to use these facilities. This work was also achieved using SQUID at the Cybermedia Center, Osaka University.
@@
  • Shiota et al. 2024 T. Shiota, K. Ishihara, T. M. Do, T. Mori and W. Mizukami, arXiv preprint arXiv:2412.13088, 2024.
@@
  • Calle-Vallejo et al. 2015 F. Calle-Vallejo, J. Tymoczko, V. Colic, Q. H. Vu, M. D. Pohl, K. Morgenstern, D. Loffreda, P. Sautet, W. Schuhmann and A. S. Bandarenka, Science, 2015, 350, 185–189.
@@
  • Roy et al. 2021 D. Roy, S. C. Mandal and B. Pathak, ACS Applied Materials & Interfaces, 2021, 13, 56151–56163.
  • Roy et al. 2024 D. Roy, S. Charan Mandal, A. Das and B. Pathak, Chemistry–A European Journal, 2024, 30, e202302679.
@@
  • Jain et al. 2013 A. Jain, S. P. Ong, G. Hautier, W. Chen, W. D. Richards, S. Dacek, S. Cholia, D. Gunter, D. Skinner, G. Ceder and K. A. Persson, APL Materials, 2013, 1, 011002.
@@
  • 90 https://github.com/Qulacs-Osaka/scikit-qulacs.